OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (g o f)(3)? 6 x + 8 -10 23 I'm so lost

OpenStudy (anonymous):

first find \(f(3)\) what do you get?

OpenStudy (anonymous):

if it is not clear let me know and i will show you

OpenStudy (anonymous):

Idk how to do it can you take me step by step

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[ f(x) = -2x + 8\] is what you are given so \[ f(3) = -2\times 3 + 8\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in other words, where you see an \(x\) you put a \(3\)

OpenStudy (anonymous):

ooooooo

OpenStudy (anonymous):

arithmetic gives \[f(3)=-2\times 3+8=-6+8=2\] so \[f(3)=2\]

OpenStudy (anonymous):

we are not done, because you were not asked for \(f(3)\) you were asked for \[g\circ f(3)=g(f(3))\] but we are done with step one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

step 2 now that we know \(f(3)=2\) what is \(g(2)\) ?

OpenStudy (anonymous):

clear or not? since \[g(x)=3x\\g(2)=3\times 2\]

OpenStudy (anonymous):

6

OpenStudy (anonymous):

that is the final answer putting it all together \[g\circ f(3)=g(f(3))=g(2)=6\]

OpenStudy (anonymous):

ooooo I feel really slow but I get it now thank you so much for your help

OpenStudy (anonymous):

yw we can do another if you like

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

you got one?

OpenStudy (anonymous):

or you want me to make one up>?

OpenStudy (anonymous):

I dont have another one like that but I have another problem

OpenStudy (anonymous):

ok lets try it

OpenStudy (anonymous):

nope, runtime error got a screenshot?

OpenStudy (anonymous):

Which of the following systems of inequalities would produce the region indicated on the graph below

OpenStudy (anonymous):

one line is \(y=x+2\) the other is \(y=-x+4\)

OpenStudy (anonymous):

your region is below both do you have an answer that looks like \[y<x+2\\ y<-x+4\]?

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

b55d5c96-0418-4ab2-9d6a-8016f2f02698.gif

OpenStudy (anonymous):